# Scalable Data Synthesis for Computer Use Agents with Step-Level Filtering

### v1

@@ Abstract
Computer use agents (CUAs) can operate real-world digital interfaces but remain difficult to train due to the high cost of graphical user interface (GUI) interaction and the scarcity of high-quality trajectory data. Existing datasets rely on human demonstrations, limiting scalability. A natural alternative is to synthesize data from strong CUAs, yet their rollouts are highly noisy, with incorrect or suboptimal actions consisting a large proportion of the steps, making naive imitation ineffective. To tackle this challenge, we introduce a scalable data synthesis pipeline that transforms noisy rollouts into reliable supervision without human annotation. The core idea is step-level filtering, which evaluates actions individually to retain only correct steps, complemented by reasoning augmentation for improved planning. Using this pipeline, we construct WebSTAR, a dataset of 13.3K trajectories and 100K graded, reasoning-rich steps synthesized from OpenAI’s computer-use-preview model. We train Qwen-2.5-VL-Instruct models (7B and 32B) on WebSTAR. On WebVoyager, our 7B model surpasses SoTA open-source CUA model UI-TARS-1.5-7B by more than 15% with only supervised finetuning. Building on step-level grading, we further create WebSCORE, a dataset of graded step-level actions, and train StepRM, a 7B multimodal reward model distilled from o4-mini, which matches its grading quality while being far more efficient to deploy at scale. Our results establish step-level filtering as a key principle for scalable CUA training, and we introduce two new datasets (WebSTAR, WebSCORE) along with a lightweight reward model (StepRM) as practical tools to advance robust and efficient CUAs. Our code is available at https://github.com/yifei-he/CUAStepFiltering.
Scalable Data Synthesis for Computer Use Agents with Step-Level Filtering
Yifei Hethanks: Work done during an internship at Microsoft. Correspondence email: yifeihe3@illinois.edu.1, Pranit Chawla2, Yaser Souri2, Subhojit Som2, Xia Song2 1University of Illinois Urbana-Champaign 2Microsoft
@@ 1 Introduction
Autonomous agents are increasingly capable of interpreting task instructions, perceiving environments, reasoning and planning, and executing precise actions. Among them, computer use agents (CUAs) (operator; Claude-computer-use; agasheagent; qin2025ui; wang2025ui; wang2025opencua) have emerged as a powerful class of models that can operate real-world digital interfaces such as web browsers (deng2023mind2web; he2024webvoyager; xue2025illusion) and desktop environments (xie2024osworld; bonattiwindows). Despite recent progress, training CUAs remains fundamentally challenging due to two key issues: the high cost of interaction with GUI environments, and the scarcity of high-quality training data.
@@
Live interaction with GUIs typically incurs high infrastructure costs (e.g., virtual machines or cloud environments (bonattiwindows)), which limits the scale at which interaction data can be collected. As a result, most CUA training is conducted offline, where models are trained on pre-collected trajectories using supervised finetuning (SFT) (xuaguvis; qin2025ui). However, offline CUA trajectory data is extremely limited. Existing data sources are either (i) manually labeled (deng2023mind2web; xuaguvis), incurring high labor costs, or (ii) curated using general-purpose vision-language models (VLMs) such as GPT-4o (he2024webvoyager; pahuja2025explorer), which are not grounded in agentic tasks and often rely on auxiliary information such as accessibility trees that are not available in general deployment settings.
A promising alternative is to synthesize trajectory data using strong CUAs themselves, such as OpenAI CUA (operator) or Claude Computer Use (Claude-computer-use). This approach resembles hard distillation in language model training, where a teacher model generates outputs and a student model learns to imitate them. However, unlike textual LLM outputs, CUA trajectories are inherently noisy. We observe that even state-of-the-art CUAs produce incorrect or suboptimal actions in more than half of the steps. Training directly on such noisy rollouts leads to degraded performance in student models, underscoring the need for more principled filtering and augmentation strategies to transform raw rollouts into high-quality training data.
@@
    Thought augmentation: Proprietary CUAs do not produce explicit reasoning traces. Thus, inspired by prior work on reasoning-augmented agents (yaoreact; xuaguvis; yang2024react), we insert an intermediate thought before each action to strengthen the agent’s reasoning and planning abilities. These thoughts provide interpretable, step-by-step guidance that grounds the action in a structured decision-making process.
@@
To evaluate the effectiveness of our data synthesis pipeline, we synthesize data with OpenAI CUA [^1] OpenAI CUA corresponds to the OpenAI computer-use-preview API 2025-03-01 release. We use Azure OpenAI for all API callling. using queries from OpenWebVoyager (he2024openwebvoyager). From these rollouts, we construct WebSTAR (WebVoyager Step-Level Trajectories with Augmented Reasoning), a dataset of 13.3K trajectories and 100K steps enriched with reasoning traces and step-level scores. To complement it, we further derive WebSCORE (WebVoyager Step-level CORrectness Evaluation), a dedicated step-level grading dataset on 50K steps for training reward models. We train Qwen-2.5-VL-Instruct-7B and 32B (Qwen2.5-VL) on WebSTAR to assess the pipeline’s effectiveness. Our contributions are summarized as follows:
@@
    WebSTAR & WebSCORE datasets: We construct WebSTAR for trajectory training with reasoning traces, and WebSCORE for reward modeling, providing high-quality resources for future CUA research.
@@
    StepRM: We further train a lightweight 7B reward model on WebSCORE, achieving grading quality comparable to o4-mini at a fraction of the cost, enabling scalable step-level filtering and paving way for RL-based optimization.
@@ 2 Preliminaries
A CUA operates by receiving a task instruction and interacting with a graphical user interface to complete the task. At each step $i$, the agent observes the current state of the interface as a screenshot $o_{i}$, and executes an action $a_{i}$ in response. Given apriori knowledge of task difficulty, a maximum number of steps $T$ is defined, within which the agent must complete the task. To enhance the agent’s reasoning and planning capabilities, it is common to adopt the ReAct framework (yaoreact), which augments each step with an intermediate thought $t_{i}$ to guide the action $a_{i}$. To ensure broad generalizability, we do not provide the agent with auxiliary inputs such as accessibility trees or specialized annotations like Set-of-Marks used in prior work (he2024webvoyager; pahuja2025explorer).
The agent is trained to generate the next thought and action $(t_{n},a_{n})$, conditioned on the task instruction $\mathcal{I}$, the sequence of past thoughts and actions $\{(t_{n-i},a_{n-i})\}_{i=1}^{n-1}$, and a truncated window of the most recent $w$ screenshots $\{o_{n-i}\}_{i=1}^{w}$. Only the image inputs are truncated due to their high token cost, which often dominates the input length in multimodal models. The action space is specified Table˜1, where we generally follow common desktop operations defined in UI-TARS (qin2025ui).
@@ 3.1 Teacher CUA Rollout
We use queries from the training split of OpenWebVoyager (he2024openwebvoyager) and collect trajectories by rolling out the OpenAI computer-use-preview API. For each query, we generate 16 rollouts with a maximum horizon of 100 steps, resulting in 21K raw trajectories across 1,326 queries. Following prior works (he2024webvoyager; xue2025illusion), we perform first-round trajectory-level filtering to remove failed rollouts by providing GPT-4o (hurst2024gpt) with the full action sequence and associated screenshots and prompting it to judge task success. After this initial grading, 13,338 successful trajectories are retained.
@@ 3.2 Step-Level Filtering
Previous attempts at data synthesis for CUAs often only retain the trajectories that successfully complete the task (murty2024bagel; pahuja2025explorer). However, the trajectory-level approach does not align well with the step-level nature of CUA training, as in Eq. 2. The training objective focuses on predicting the next thought and action at each individual step, rather than modeling the entire trajectory end-to-end. The key limitation of trajectory-level filtering is that even successful trajectories often contain suboptimal or incorrect intermediate steps. Directly using these full trajectories for SFT exposes the model to flawed decision points, which can bias it toward replicating incorrect behaviors and ultimately degrade downstream performance.
@@
To address the limitations of trajectory-level filtering, we propose a step-level filtering strategy that further evaluates each individual action in the successful trajectories. Instead of retaining entire trajectories, our method assigns a quality score to each step based on its expected contribution to task success. This evaluation is performed by a grading model that reasons over the context including the task instruction, the current and past actions, a window of the past $w$ screenshots annotated with actions following chen2024guicourse, and a zoomed-in crop of the current screenshot centered on the action target, illustrated in Figure˜1. We use o4-mini as our grading model due to its strong reasoning capability and relatively lightweight inference cost. Importantly, o4-mini provides structured intermediate analysis, which enables more consistent judgments across diverse task types. The grading process proceeds in four stages:
@@ 3.3 Thought Augmentation
The raw trajectories collected from OpenAI CUA consist solely of sequences of observations and actions, without any explicit reasoning traces. However, prior work has shown that intermediate reasoning steps are crucial for enhancing agents’ planning and decision-making capabilities, especially in complex, multi-step tasks (yaoreact).
To address this, we follow previous approaches (xuaguvis; yang2024react; qin2025ui) and generate thoughts in a post-hoc manner. Each thought is designed to bridge perception and action, and is structured around three key components: i) Situation description: A brief description of the current observation, especially focusing on the region where the action takes place. ii) Reasoning: An explanation of the rationale behind the action and how it contributes to completing the overall task. iii) Actionable instruction: A natural language command that describes the action in an interpretable way, serving as a textual bridge between thought and execution. The detailed prompt is presented in Appendix˜F.
@@
The results in Table 3 show that incorporating thought augmentation yields consistent benefits, especially under the Pass@4 metric. In line with prior findings (xuaguvis), the Pass@1 scores with and without thought traces remain close, which is expected since Pass@1 is a brittle indicator of reliability: CUA policies are stochastic, and in long-horizon tasks, a single error can derail an entire trajectory. By contrast, Pass@4 provides a more faithful measure of performance, reflecting whether the agent can succeed across multiple sampled rollouts rather than relying on a single attempt.
@@ 4.1 Setup
We roll out OpenAI CUA on queries from the OpenWebVoyager (he2024openwebvoyager) dataset. The original query set comes from the training split of both WebVoyager (he2024webvoyager) and Mind2Web (deng2023mind2web), totaling 2,111 queries. However, due to the dynamic nature of websites, many tasks become outdated or invalid, caused by issues such as websites access errors, added human verification steps, or updated URLs. After filtering out these queries, we retain 1,326 valid ones. Details about those website failures are discussed in Appendix˜B.
To interact with live websites, we adopt the WebVoyager (he2024webvoyager) codebase. However, the original implementation only supports Set-of-Mark (SoM) prompting, which is incompatible with the more common raw coordinate interaction. Thus, we modify the implementation with Playwright to enable more robust interactions with websites. For each query, we cap the rollout length at 100 steps (i.e., $T=100$), terminating the episode at step 100 regardless of whether the task is completed.
@@
We use Qwen-2.5-VL-7B-Instruct and Qwen-2.5-VL-32B-Instruct (Qwen2.5-VL) as base models. Following common practice (wang2024qwen2; xuaguvis), we freeze the vision tower and only train the language components. All experiments are conducted on NVIDIA A100 GPUs with 80 GB memory, using LlamaFactory (zheng2024llamafactory) for multimodal SFT. We set the sequence length to 8,192 and feed in only the most recent screenshot ($w=1$) following xuaguvis, as this configuration yields the most stable training in our setting. We use an initial learning rate of 1e-5 with warmup ratio of 0.03 and cosine decay to 0. The global batch size is set to 128. Training on WebSTAR for 2 epochs takes approximately 256 GPU hours for 7B models and 1024 GPU hours for 32B models.
@@
To test the capabilities of CUA, we deploy three online evaluation benchmarks: WebVoyager (he2024webvoyager), Mind2Web-Live (pan2024webcanvas) and Online Mind2Web (xue2025illusion). The online evaluation is particularly challenging due to the everchanging nature of websites, requiring the model to generalize the static data during offline training to dynamically changing environments. Following he2024webvoyager, we provide the full action trajectory and corresponding screenshots of the agent to GPT-4o (hurst2024gpt) to automatically judge the task success.
@@ 4.4 StepRM
To alleviate this bottleneck, we introduce StepRM, a lightweight multimodal reward model based on Qwen2.5-VL-7B. StepRM is explicitly trained to replicate the step-level grading behavior of o4-mini, effectively distilling its judgments into a compact, open model. The training is conducted on the WebSCORE dataset, which contains 200K graded steps randomly sampled from the raw rollout trajectories collected in our data synthesis pipeline. To ensure balanced supervision, the dataset contains 50% steps with score >5>5 and 50% with score $\leq 5$. Each sample includes the same multimodal inputs used in our grading pipeline (the latest annotated screenshot, recent history of actions, and task instruction) and predicts a scalar quality score for the current action. In this way, StepRM functions as a direct substitute for o4-mini in step-level filtering, while being significantly cheaper to deploy at scale. StepRMuses the same hyperparameters and compute setup as our main policy models.
Table 7: Step-level filtering performance from StepRM with Pass@1 (Pass@4). Compared with o4-mini as a grading model, StepRM achieves competitive quality while being substantially more efficient, demonstrating that lightweight learned reward models can replace expensive LLM evaluations for step-level filtering.
@@
Beyond serving as a cost-effective substitute for step-level filtering, StepRM also has the potential to act as a critic in reinforcement learning frameworks such as PPO (schulman2017proximal). In this setting, StepRM can provide dense, step-level feedback to the agent, helping to overcome two central challenges in CUA training: sparse rewards, where task success is only known at the end of long trajectories, and credit assignment, where it is difficult to determine which intermediate actions led to success or failure. By offering fine-grained step evaluations, StepRM could enable more sample-efficient RL training of CUAs in interactive environments. We leave the integration of StepRM into RL pipelines as an exciting direction for future work.

### v2

# WebSTAR: Scalable Data Synthesis for Computer Use Agents with Step-Level Filtering

@@ Abstract
Computer use agents (CUAs) can operate real-world digital interfaces but remain difficult to train due to the high cost of graphical user interface (GUI) interaction and the scarcity of high-quality trajectory data. Existing datasets rely on human demonstrations, limiting scalability. A natural alternative is to synthesize data from strong CUAs, yet their rollouts are highly noisy, with incorrect or suboptimal actions consisting a large proportion of the steps, making naive imitation ineffective. To tackle this challenge, we introduce a scalable data synthesis pipeline that transforms noisy rollouts into reliable supervision without human annotation. The core idea is step-level filtering, which evaluates actions individually to retain only correct steps, complemented by reasoning augmentation for improved planning. Using this pipeline, we construct WebSTAR, a dataset of 13.3K trajectories and 100K graded, reasoning-rich steps synthesized from OpenAI’s computer-use-preview model. We train Qwen-2.5-VL-Instruct models (7B and 32B) on WebSTAR. On WebVoyager, our 7B model surpasses SoTA open-source CUA model UI-TARS-1.5-7B by more than 15% with only supervised finetuning. Building on step-level grading, we further create WebSCORE, a dataset of graded step-level actions, and train StepRM, a 7B multimodal process reward model distilled from o4-mini, which matches its grading quality while being far more efficient to deploy at scale. Our results establish step-level filtering as a key principle for scalable CUA training, and we introduce two new datasets (WebSTAR, WebSCORE) along with a lightweight process reward model (StepRM) as practical tools to advance robust and efficient CUAs.
WebSTAR: Scalable Data Synthesis for Computer Use Agents with Step-Level Filtering
Yifei Hethanks: Work done during an internship at Microsoft. Correspondence email: yifeihe3@illinois.edu.1, Pranit Chawla2, Yaser Souri2, Subhojit Som2, Xia Song2 1University of Illinois Urbana-Champaign 2Microsoft Website[Uncaptioned image] Dataset Code
@@ 1 Introduction
Autonomous agents are increasingly capable of interpreting task instructions, perceiving environments, reasoning and planning, and executing precise actions. Among them, computer use agents (CUAs) (OpenAI, 2025; Anthropic, 2024; Agashe et al., 2025; Qin et al., 2025; Wang et al., 2025a, b) have emerged as a powerful class of models that can operate real-world digital interfaces such as web browsers (Deng et al., 2023; He et al., 2024a; Xue et al., 2025) and desktop environments (Xie et al., 2024; Bonatti et al., 2025; Wu et al., 2025). Despite recent progress, training CUAs remains fundamentally challenging due to two key issues: the high cost of interaction with GUI environments, and the scarcity of high-quality training data.
@@
Live interaction with GUIs typically incurs high infrastructure costs (e.g., virtual machines or cloud environments (Bonatti et al., 2025)), which limits the scale at which interaction data can be collected. As a result, most CUA training is conducted offline, where models are trained on pre-collected trajectories using supervised finetuning (SFT) (Xu et al., 2025b; Qin et al., 2025). However, offline CUA trajectory data is extremely limited. Existing data sources are either (i) manually labeled (Deng et al., 2023; Xu et al., 2025b), incurring high labor costs, or (ii) curated using general-purpose vision-language models (VLMs) such as GPT-4o (He et al., 2024a; Pahuja et al., 2025), which are not grounded in agentic tasks and often rely on auxiliary information such as accessibility trees that are not available in general deployment settings.
A promising alternative is to synthesize trajectory data using strong CUAs themselves, such as OpenAI CUA (OpenAI, 2025) or Claude Computer Use (Anthropic, 2024). This approach resembles hard distillation in language model training, where a teacher model generates outputs and a student model learns to imitate them. However, unlike textual LLM outputs, CUA trajectories are inherently noisy. We observe that even state-of-the-art CUAs produce incorrect or suboptimal actions in more than half of the steps. Training directly on such noisy rollouts leads to degraded performance in student models, underscoring the need for more principled filtering and augmentation strategies to transform raw rollouts into high-quality training data.
@@
    Thought augmentation: Proprietary CUAs do not produce explicit reasoning traces. Thus, inspired by prior work on reasoning-augmented agents (Yao et al., 2023; Xu et al., 2025b; Yang et al., 2024), we insert an intermediate thought before each action to strengthen the agent’s reasoning and planning abilities. These thoughts provide interpretable, step-by-step guidance that grounds the action in a structured decision-making process.
@@
To evaluate the effectiveness of our data synthesis pipeline, we synthesize data with OpenAI CUA [^1] OpenAI CUA corresponds to the OpenAI computer-use-preview API 2025-03-01 release. We use Azure OpenAI for all API callling. using queries from OpenWebVoyager (He et al., 2024b). From these rollouts, we construct WebSTAR (WebVoyager Step-Level Trajectories with Augmented Reasoning), a dataset of 13.3K trajectories and 100K steps enriched with reasoning traces and step-level scores. To complement it, we further derive WebSCORE (WebVoyager Step-level CORrectness Evaluation), a dedicated step-level grading dataset on 50K steps for training process reward models. We train Qwen-2.5-VL-Instruct-7B and 32B (Bai et al., 2025b) on WebSTAR to assess the pipeline’s effectiveness. Our contributions are summarized as follows:
@@
    WebSTAR & WebSCORE datasets: We construct WebSTAR for trajectory training with reasoning traces, and WebSCORE for process reward modeling, providing high-quality resources for future CUA research.
@@
    StepRM: We further train a lightweight 7B process reward model on WebSCORE, achieving grading quality comparable to o4-mini at a fraction of the cost, enabling scalable step-level filtering and paving way for RL-based optimization.
@@ 2 Preliminaries
A CUA operates by receiving a task instruction and interacting with a graphical user interface to complete the task. At each step $i$, the agent observes the current state of the interface as a screenshot $o_{i}$, and executes an action $a_{i}$ in response. Given apriori knowledge of task difficulty, a maximum number of steps $T$ is defined, within which the agent must complete the task. To enhance the agent’s reasoning and planning capabilities, it is common to adopt the ReAct framework (Yao et al., 2023), which augments each step with an intermediate thought $t_{i}$ to guide the action $a_{i}$. To ensure broad generalizability, we do not provide the agent with auxiliary inputs such as accessibility trees or specialized annotations like Set-of-Marks used in prior work (He et al., 2024a; Pahuja et al., 2025).
The agent is trained to generate the next thought and action $(t_{n},a_{n})$, conditioned on the task instruction $\mathcal{I}$, the sequence of past thoughts and actions $\{(t_{n-i},a_{n-i})\}_{i=1}^{n-1}$, and a truncated window of the most recent $w$ screenshots $\{o_{n-i}\}_{i=1}^{w}$. Only the image inputs are truncated due to their high token cost, which often dominates the input length in multimodal models. The action space is specified Table˜1, where we generally follow common desktop operations defined in UI-TARS (Qin et al., 2025).
@@ 3.1 Teacher CUA Rollout
We use queries from the training split of OpenWebVoyager (He et al., 2024b) and collect trajectories by rolling out the OpenAI computer-use-preview API. For each query, we generate 16 rollouts with a maximum horizon of 100 steps, resulting in 21K raw trajectories across 1,326 queries. Following prior works (He et al., 2024a; Xue et al., 2025), we perform first-round trajectory-level filtering to remove failed rollouts by providing GPT-4o (Hurst et al., 2024) with the full action sequence and associated screenshots and prompting it to judge task success. After this initial grading, 13,338 successful trajectories are retained.
@@ 3.2 Step-Level Filtering
Previous attempts at data synthesis for CUAs often only retain the trajectories that successfully complete the task (Murty et al., 2024; Pahuja et al., 2025). However, the trajectory-level approach does not align well with the step-level nature of CUA training, as in Eq. 2. The training objective focuses on predicting the next thought and action at each individual step, rather than modeling the entire trajectory end-to-end. The key limitation of trajectory-level filtering is that even successful trajectories often contain suboptimal or incorrect intermediate steps. Directly using these full trajectories for SFT exposes the model to flawed decision points, which can bias it toward replicating incorrect behaviors and ultimately degrade downstream performance.
@@
To address the limitations of trajectory-level filtering, we propose a step-level filtering strategy that further evaluates each individual action in the successful trajectories. Instead of retaining entire trajectories, our method assigns a quality score to each step based on its expected contribution to task success. This evaluation is performed by a grading model that reasons over the context including the task instruction, the current and past actions, a window of the past $w$ screenshots annotated with actions following Chen et al. (2024), and a zoomed-in crop of the current screenshot centered on the action target, illustrated in Figure˜1. We use o4-mini as our grading model due to its strong reasoning capability and relatively lightweight inference cost. Importantly, o4-mini provides structured intermediate analysis, which enables more consistent judgments across diverse task types. The grading process proceeds in four stages:
@@ 3.3 Thought Augmentation
The raw trajectories collected from OpenAI CUA consist solely of sequences of observations and actions, without any explicit reasoning traces. However, prior work has shown that intermediate reasoning steps are crucial for enhancing agents’ planning and decision-making capabilities, especially in complex, multi-step tasks (Yao et al., 2023).
To address this, we follow previous approaches (Xu et al., 2025b; Yang et al., 2024; Qin et al., 2025) and generate thoughts in a post-hoc manner. Each thought is designed to bridge perception and action, and is structured around three key components: i) Situation description: A brief description of the current observation, especially focusing on the region where the action takes place. ii) Reasoning: An explanation of the rationale behind the action and how it contributes to completing the overall task. iii) Actionable instruction: A natural language command that describes the action in an interpretable way, serving as a textual bridge between thought and execution. The detailed prompt is presented in Appendix˜F.
@@
The results in Table 3 show that incorporating thought augmentation yields consistent benefits, especially under the Pass@4 metric. In line with prior findings (Xu et al., 2025b), the Pass@1 scores with and without thought traces remain close, which is expected since Pass@1 is a brittle indicator of reliability: CUA policies are stochastic, and in long-horizon tasks, a single error can derail an entire trajectory. By contrast, Pass@4 provides a more faithful measure of performance, reflecting whether the agent can succeed across multiple sampled rollouts rather than relying on a single attempt.
@@ 4.1 Setup
We roll out OpenAI CUA on queries from the OpenWebVoyager (He et al., 2024b) dataset. The original query set comes from the training split of both WebVoyager (He et al., 2024a) and Mind2Web (Deng et al., 2023), totaling 2,111 queries. However, due to the dynamic nature of websites, many tasks become outdated or invalid, caused by issues such as websites access errors, added human verification steps, or updated URLs. After filtering out these queries, we retain 1,326 valid ones. Details about those website failures are discussed in Appendix˜B.
To interact with live websites, we adopt the WebVoyager (He et al., 2024a) codebase. However, the original implementation only supports Set-of-Mark (SoM) prompting, which is incompatible with the more common raw coordinate interaction. Thus, we modify the implementation with Playwright to enable more robust interactions with websites. For each query, we cap the rollout length at 100 steps (i.e., $T=100$), terminating the episode at step 100 regardless of whether the task is completed.
@@
We use Qwen-2.5-VL-7B-Instruct and Qwen-2.5-VL-32B-Instruct (Bai et al., 2025b) as base models. Following common practice (Wang et al., 2024; Xu et al., 2025b), we freeze the vision tower and only train the language components. All experiments are conducted on NVIDIA A100 GPUs with 80 GB memory, using LlamaFactory (Zheng et al., 2024b) for multimodal SFT. We set the sequence length to 8,192 and feed in only the most recent screenshot ($w=1$) following Xu et al. (2025b), as this configuration yields the most stable training in our setting. We use an initial learning rate of 1e-5 with warmup ratio of 0.03 and cosine decay to 0. The global batch size is set to 128. Training on WebSTAR for 2 epochs takes approximately 256 GPU hours for 7B models and 1024 GPU hours for 32B models.
@@
To test the capabilities of CUA, we deploy three online evaluation benchmarks: WebVoyager (He et al., 2024a), Mind2Web-Live (Pan et al., 2024) and Online Mind2Web (Xue et al., 2025). The online evaluation is particularly challenging due to the everchanging nature of websites, requiring the model to generalize the static data during offline training to dynamically changing environments. Following He et al. (2024a), we provide the full action trajectory and corresponding screenshots of the agent to GPT-4o (Hurst et al., 2024) to automatically judge the task success.
@@ 4.4 StepRM
To alleviate this bottleneck, we introduce StepRM, a lightweight multimodal process reward model based on Qwen2.5-VL-7B. StepRM is explicitly trained to replicate the step-level grading behavior of o4-mini, effectively distilling its judgments into a compact, open model. The training is conducted on the WebSCORE dataset, which contains 200K graded steps randomly sampled from the raw rollout trajectories collected in our data synthesis pipeline. To ensure balanced supervision, the dataset contains 50% steps with score >5>5 and 50% with score $\leq 5$. Each sample includes the same multimodal inputs used in our grading pipeline (the latest annotated screenshot, recent history of actions, and task instruction) and predicts a scalar quality score for the current action. In this way, StepRM functions as a direct substitute for o4-mini in step-level filtering, while being significantly cheaper to deploy at scale. StepRMuses the same hyperparameters and compute setup as our main policy models.
Table 7: Step-level filtering performance from StepRM with Pass@1 (Pass@4). Compared with o4-mini as a grading model, StepRM achieves competitive quality while being substantially more efficient, demonstrating that lightweight learned process reward models can replace expensive LLM evaluations for step-level filtering.
@@
Beyond serving as a cost-effective substitute for step-level filtering, StepRM also has the potential to act as a critic in reinforcement learning frameworks such as PPO (Schulman et al., 2017). In this setting, StepRM can provide dense, step-level feedback to the agent, helping to overcome two central challenges in CUA training: sparse rewards, where task success is only known at the end of long trajectories, and credit assignment, where it is difficult to determine which intermediate actions led to success or failure. By offering fine-grained step evaluations, StepRM could enable more sample-efficient RL training of CUAs in interactive environments. We leave the integration of StepRM into RL pipelines as an exciting direction for future work.